Campus Technology Services

Taking Steps for the Future

TechNews - April 2013

This is a summary of the activities supporting faculty, staff and student technology services at SUNY Oswego that occurred during March 2013.

Reminders - Please make sure that your co-workers know about these

  1. Routine Service Outage
    As part of Campus Technology Services' efforts to keep servers running at optimum levels, a regular maintenance period has been designated. From 6AM to 10AM on the second Sunday of each month, several services such as E-mail, editing of both departmental and personal web pageswill not be available. The next scheduled outage is Sunday, April 14th, 2013.
  2. NYALERT
    For information on NYAlert: http://www.oswego.edu/administration/public_affairs/emergency/nyalert.html or 
    visit http://www.suny.edu/SUNYALERTsec to sign up.
  3. RAVE Guardian
    For information on Rave Guardian http://oswego.edu/student/blogs/?cat=11 or to sign up: https://www.getrave.com/login/oswego
  4. Patch Wednesdays
    CTS now has the capability of updating your campus owned computers for security fixes as well as functionality on a regularly scheduled basis. This may include software updates to programs such as Adobe Reader or Microsoft Word or to change Windows settings to better support your computers. In order to streamline this process we will only perform these updates on Wednesdays.  These changes will be implemented only after thorough internal CTS testing. You may possibly notice a slight slowdown in computer performance at that time. Updates are added to a CTS calendar so you may see what changes have occurred to your computer on any given day or that are scheduled to be done.
  5. Workshops calendar 
    Look here for a listing of all professional development workshop registrations, calendars and information. 
  6. CTS News
    Make sure to visit our CTS New blog for some of our latest information on updated or new services, future services, technology tips, and more.
  7. CTS Projects
    Check our CTS Projects for brief descriptions of the projects CTS is currently working on. 

Technology Support Center

TSC Staff have been working on implementing the new Service Now ticketing system.  We will be replacing the HEAT ticketing system with the more robust, easy to use, ITIL-based ServiceNow product. Our first round of training with CTS staff has been completed.  We will be working on improvements based on feedback and then do another round of training.

We will be searching for a new Help Desk Support Specialist full-time position.  HR is working on advertising the position.

We completed our student recruitment efforts and will now be reviewing applications to hire new Help Desk Assistants and Computer Lab Assistants for the 2013-2014 academic year.

You can reach us by phone at 312.3456 or on site at 26 Lanigan Hall, on Facebook at "SUNY Oswego - Campus Technology Services" on Twitter at www.twitter.com/sunyoswegocts or our website.

Learning Management System

Please check the ANGEL website for dates and times for upcoming workshops.

The Blackboard migration timeline has been finalized with the SUNY Learning Network.  We will be looking at doing a pilot during the Spring 2015 semester and then move into full production during summer 2015.

If you are interested in setting up a community group, please visit the ANGEL web site at www.oswego.edu/academics/angel. Community groups are limited to academic and/or university related activities only.

Professional Development

Spring workshops are winding down.    There will be a few more available the first couple of weeks in April.  They can be viewed from our CTS Workshops Registration page.

If you missed the Cyber Security Awareness trainings offered last fall and in January, you may view a recording of it on our CTS Security site.  You will need to enter your Laker NetID to view it.

There is a new professional development website that includes links to training schedules and registrations for workshops offered by CELT, ORSP, ANGEL, Web Development (Ingeniux) and CTS.  Check it out at www.oswego.edu/workshops.

Self-paced, on-demand tutorials are also available from our iTunes U site.  Tutorials covering Microsoft Word 2010, Excel 2010, PowerPoint 2010 and LakerApps Sites are currently available.

The Skillport e-learning software is available to all faculty, staff and students.  Skillport includes business and technology-related courses (including Office 2010) as well as courses and test preps needed to prepare you for many business and technical certifications.   Get started at the Skillport home page.

Software News

SUNY Oswego faculty and staff may purchase a copy of either MS Office Professional Plus 2010 (PC version) or Office 2011 (Mac version) for use on their personal computers for $9.95 plus tax.  This "Home Use Program" is made available to faculty and staff through the campus purchase of the Microsoft Campus Agreement.  Details of how to go about purchasing can be found from the CTS MS Office Home Use Program (HUP) page.

Microsoft offers students a significantly reduced price on its products.  For the special pricing and to buy online visit www.theultimatesteal.com.

Web Update

Campus Technology Services (CTS) would like to inform users that new releases of browsers aren't always fully compatible with some web-based campus services that you need to access.  The use of an unsupported browser can result in inconsistent or unpredictable behavior when working in applications.  Please check with the TSC before updating.

 

Desktop Support

During the month of March, Campus Technology Services technicians have continued to aid faculty, staff and students in resolving hardware, software, and account related issues. Since the beginning of March Campus Technology Services technicians have assisted in the resolution of 51 hardware, 60 software, and 27 account related support requests.

Equipment Deployments

Our desktop support technicians continue to prepare and deploy new equipment as it arrives. During the month of March new equipment has been configured and installed for: Campus Technology Services, Chemistry Department, Computer Science Department, Earth Sciences Department, Education Administration Department, History Department, Interdisciplinary Programs and Activities Center, Philosophy Department, Psychology Department, Registrar's Office, and School of Business.

Lab Utilization Statistics

The table below displays the number of login sessions that have occurred within the campus open access computer labs over the last month.

Location # of Logins
CC_207 4037
Hart_Lab 746
Mahar_210 615
Mahar_211 900
Oneida_Lab 571
Penfield_24_Hour 2439
Penfield_Lab 5407
Penfield_Laptops 606
Penfield_Lobby 3931
Scales_Lab 317
Snygg_6 764
VDI 528

During the month of March, 59 virtual applications were launched.

Banner

  • Admissions - Developed a process to update (aged) recruit (only) data so that the people will not be included in common matching of new data; resolvedd ASC load problems along with SICAS, PSAT recruitload review, modified  four programs as requested
  • Advancement/Development - developed programs for gift analysis, researched and outlined the use of AGTOGIV to be used by Web Services, trained staff to use the SSB web product for prospecting
  • Athletics - Developed an (argos) extract of athletes for ASC Athletics implementation; set up the midterm grade emails
  • Financial Aid - set up and tested 1314 application loads, set up and executed scholarship rolls for new admits and returning students, set up batch posting for IRS tax return matches to comply with new Fed standards, adapted award and MIL letters, developed an overpaid loan report
  • Housing - Assisted Res Life & Housing in set up of Banner rules and Adirondack Returning Student lotteries, giving extra attention to Spring admits
  • Institutional Research - SIRIS support, corrected (A14018) processing errors as requested, modified five programs as requested
  • Library - modified (A57199) to identify students attending classes at the Metro campus
  • Parking - AIMS support (prorate schedule, documentation)
  • Provost (Extended Learning) - beta tested the Veteran Survey application developed by SICAS, participated in discussions for selection of an Early Alert product (Starfish Early Alert), participated in the Gates Foundation Grant application
  • Registration - Degreeworks implemenation (access, FAQs, business process definition), updated the non-degree WebApp process to require Registrar review prior to enrollment, modified five programs as requested
  • Student Advisement - New Student Menu adapted to determine if the student is transfer, freshmen, or non-degree; corrected housing link; Medicat implementation review; assisted with new information sheet PDF process and later developed a program to assist advisement with indexing information sheets, modified P21020 to conform with new business processes, Vizual Zen support, no longer adding a checklist item (P90007) for math placement exams, reviewed and set up Faculty Feedback on bantest

Other Administrative Support

  • Banner application server upgrade/migration completed
  • Banner Upgrade - installed several Ellucian General and Student Releases in the production database.  This included adding local modifications to baseline objects (.sqr, .pc. fmb)
  • Banner Patches - reviewed list of uninstalled patches to identify which patches needed application after the upgrade (above), installed on test and prod databases
  • CAS Server planning with IT
  • Pinnacle upgrade planning and testing
  • BDMS configuration of reports manager on test bdms syste• Testing Windows 8
  • Upgrade HP Web Jetadmin server to version 10.3
  • Replaced landesk-console with landesk-con.
  • Updating servers to Sophos 10.2
  • Working to implement handheld scanning for assets.
  • Assisting in VDI project.

 

Provided AV support for the following on campus events  

  •  Lilly Ledbetter lecture
  • Ernst and Young presentation: Ina Issa 
  • Sportsmanship Day 

Campus Projects  

  • Shut down all ATC equipment for campus power outage and then brought classrooms back on line. 
  • Began pulling A/V cables for teaching labs in Shineman Science Building and programming those control systems
  • Design and planning for Park Hall and SEI conference rooms
  • Researched and evaluated five high lumen projectors for new large campus lecture halls
  • Designed QAM channel digital distribution for digital signage of CATV channel 1.1 for feeds to new SEI and Park Hall. 
  • CKWS now on digital 12.1

Department Support

  • Performed maintenance on WNYO transmitter and exciter
  • Installed three flat panel displays in Rich Hall
  • Demonstrated portable assisted listening system for ADA consideration
  • Review of A/V support needs for Quest
  • Provided upgrades and preventive maintenance in Metro Center classrooms
  • Provided orientation to student response systems for Provost Office
  • Installed flat panel and equipment in Penfield conference room

CTS Projects or Support

  • Upgraded Crestron control systems in nine Campus Center classrooms.
  • Installed new document cameras in CAmpus Center 211 and 306 for evaluation purposes

Telecommunications 

  • VoIP phone installations continue
  • SEI/ SOE Technology Planning/ Field Verifications
  • Fiber work for Romney Fire Alarm
  • Moves Add Changes
    • 2 Adds
    • 0 Change
    • 16 Disconnects
    • 1 Moves
  • Monthly Billing

Networking Update

  • Installed new switches, UPS and PDU units in Romney
  • Replaced King UPS unit with zero down-time after correct electrical circuit was installed, removed old unit
  • Funnelle BDF UPS failure, spent a lot of time troubleshooting with APC, they finally sent a replacement unit
  • Replaced faulty battery module in Penfield 1st floor UPS
  • Finished adding devices into Structureware Central for UPS/PDU management/monitoring and enabling centralized authentication for the devices
  • Spoke with network group at SUNY Purchase about our Cisco wireless implementation
  • Worked on 360Systems server for Pat Moochler that we could not get the ftp service to start on.  Ended up breaking into the system from a bootable CD and manually creating a root account in order to modify the ProFTPD configuration on the box, it is working now
  • Fixed Fajita Grill VPN connection for Auxiliary Services, Mark started testing a new register at this location to see if this will fix the continued issues there
  • Looked at reported wireless issue in Waterbury 2nd floor, see nothing with our equipment, client side issue
  • Modified WTOP VPN secrets after they reported some of their systems being changed without their knowledge
  • TimeWarner business class cable outage at Metro Center, fixed itself while working with Gwen
  • Created firewall rules for Ingeniux to access internal LDAP read servers for Rick Buck testing updates
  • Created firewall rules to allow LAB computers access to LDMS console
  • Created firewall rules to allow public access to SAML SSO services
  • Created firewall rules to allow public access to new MyOswego SSO login application
  • Worked on test configurations for internal routers in ATC podiums for ATC device management.  Ordered units for Poucher/Swetman upgrades and pre-configured these for the ATC group to physically install in each podium.
  • Campus wide power outage - lasted much longer than anticipated (twice as long), had to come in to shut down additional systems in order to prevent Lanigan from losing battery power.  Wireless controllers that did lose power not all come back up properly after the outage.  Had to re-sync configuration from PI to ACAD controllers that lost power in Lee to fix the problem.  Also a handful of devices did not come back with their correct naming convention, had to manually change all of these.
  • Site-fault UPS errors after power outage in Laker and several other locations on campus, opened ticket with electricians to investigate
  • Fixed access point in Hart that would not power up and corrected switchport description
  • Installed redundant fiber uplinks in Funnelle, picked up old Nortel phones that someone left in the BDF
  • Finally got all UPS units moved into their correct locations in Piez and the addition after several attempts to move (elevator wasn't working or the corridors were being worked on), still no power to any of the data closets in either section
  • Moved switches and racked them in the 1st, 2nd, 3rd floors of Piez.  Wrapped in plastic to protect from continued construction dust and debris.  Basement aggregation closet still does not have any data cables ran to it so we are unable to rack equipment there until that is complete
  • Installed updated license key on Exinda application after purchase order went through
  • Fixed Swetman wireless map dimensions
  • Troubleshoot Wireless in IPAC area in CC - user training on Cell service vs Wifi in meeting area
  • Modified temporary SAN storage for Cisco UCS box, replaced 2 drives that were failing, fixed the ZFS pool after inserting 2 SSD drives caused drive mapping to get out of whack
  • Fixed radius logs after running out of space on radius-02 over the weekend.  testlabswitch-bdf did not have its config saved before power outage
  • Spoke with Paul about Planetarium computers, need more info that Scott Steiger was suppose to forward
  • Spoke with Chuck about not being able to connect to the DMPS on 1 CC device and an additional ip/port he will need forwarded, ports are being forwarded, default gateway must be configured on his devices to work properly even though they are located on a single internal private subnet and aren't being routed anywhere.
  • XBOX/PS3/Wii/other registrations:  4
  • DMCA take down notifications:  1
  • Abuse Notifications:  3

Staff Updates

  • Apple Workshop on Deploying Apple equipment in an educational environment at Syracuse University - Andy Goldzweig, David Kahn
  • Apple Hardware Repair Certification - David Kahn
  • Internal Control Training and Quiz:  Hageny, McQuin
  • InfoComm CTS recertification:  Hageny
  • Service Now training: Canning, Hageny, McQuin, Norman, Schleinitz,
  • Western NY Cisco UCS Users Group Meeting:  Greg Fuller

Phone Calls to the Technology Support Center (3/1/13-3/31/13): 617

Call Logging Statistics - 3/1- 3/31/2013

Tickets Logged By Time Number of Tickets
8:00 AM - 11:59 AM 232
12:00 PM - 4:30 PM   219
4:31 PM - 9:00 PM 28
Off Hours

6

Total Tickets Logged 485